OpenStudy (anonymous):

Solve for x: 1 over 3 (2x − 8) = 4

OpenStudy (anonymous):

(1/3)(2x-8)=4 (2/3)x-(8/3)=4 add 8/3 (2/3)x=6 2/3 divide by 2/3 or multiply by 3/2 (still gat same answer) x=10

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

i got x=28/6 too

OpenStudy (anonymous):

you might of done something wrong @UltimateTopSpeed @nonamenever because x=10 try plugging in 28/6 into the equation and see if you get 4

OpenStudy (anonymous):

@UltimateTopSpeed, is totally right!

OpenStudy (anonymous):

on my mutiple choice it does not have 28/6

OpenStudy (anonymous):

but it did have 10

OpenStudy (anonymous):

b.c you can simplify that fraction

OpenStudy (anonymous):

how would you rediuce 28/6

OpenStudy (anonymous):

by2

OpenStudy (anonymous):

it had 10 over 3 as a choice 2

OpenStudy (anonymous):

it should be just 10

OpenStudy (anonymous):

however if you plug in 10 in the equation, that doesnt seem to b right..it ends up being 1/36=4

OpenStudy (anonymous):

dats not a choice in it tho

OpenStudy (anonymous):

@lulu22, try to reduce the 28/6, is that reduced fraction one of the choices?

OpenStudy (anonymous):

what are all the answer choices?

OpenStudy (anonymous):

no because (1/3)*(2(10)-8)=4 (1/3)*(20-8)=4 (1/3)*12=4 4=4

OpenStudy (anonymous):

i submitted the test he was right it is 10 i got it right

OpenStudy (anonymous):